# New methods for drug synergy prediction

### v1

@@ 1 Introduction
In this mini-review we focus on the new prediction methods based on deep learning and machine learning developed in last years. We put these papers under a unifying lens by first highlighting the prediction setups, included the prediction scenarios, the data sources, the input data types and synergy scores used in model output, as well as the evaluation protocols used. Following that, we discuss the core technologies, commonly used in the models, and present an overview of the predictive performance of the models.
@@ 4 Prediction performance of the models
Figure 1 summarizes the predictive performance of the models reviewed here. Figure 1a depicts the reported AUROC values for each classification method. The highest values are due to GAECDS (0.98), MatchMaker (0.97) and Kim et al. (0.96) while nine further methods reach 0.9 AUROC or more. Methods that report performance over multiple scenarios exhibit a range of values given by the Box-Whisker plot. Among regression models predicting Loewe synergy (Figure1b), MARSY obtains the highest Pearson Correlation Coefficient (PCC) of 0.89 on DrugComb, significantly above the second highest value 0.83 by MGAE-DC on Merck-2016, both of these obtained in the LTO scenario. Overall, the regression performance varies significantly between different methods. Figures 1c and 1d depict the distributions of all AUROC and PCC values, respectively, per scenario. The varying degrees of difficulty of the scenarios can be seen clearly, LTO and LPO being the easiest to predict, LDO being the most difficult and LCO the second most difficult. Figure 1e shows the reported AUROC values per dataset. The distributions are relatively similar, showing that the effect of the dataset on the results is rather moderate.

### v2

@@ 1 Introduction
In this mini-review we focus on the new prediction methods based on deep learning, pioneered by the DeepSynergy method [6], and machine learning developed in last years. We put these papers under a unifying lens by first highlighting the prediction setups, included the prediction scenarios, the data sources, the input data types and synergy scores used in model output, as well as the evaluation protocols used. Following that, we discuss the core technologies, commonly used in the models, and present an overview of the predictive performance of the models.
@@ 4 Prediction performance of the models
Figure 1 summarizes the predictive performance of the models reviewed here. Figure 1a depicts the reported AUROC values for each classification method. The highest values are due to GAECDS (0.98), MatchMaker (0.97) and Kim et al. (0.96) while nine further methods reach 0.9 AUROC or more. Methods that report performance over multiple scenarios exhibit a range of values given by the Box-Whisker plot. Among regression models predicting Loewe synergy (Figure1b), MARSY obtains the highest Pearson Correlation Coefficient (PCC) of 0.89 on DrugComb, above the second highest value 0.83 by MGAE-DC on Merck-2016, both of these obtained in the LTO scenario. Figures 1c and 1d depict the distributions of AUROC and PCC values in different scenarios. The LCO and LDO scenarios show as significantly harder prediction tasks than LPO and LTO: based on Wilcoxon rank-sum test with Bonferroni correction, LDO has significantly lower AUROC compared to LPO (p=0.026), and LTO (p=0.002). Similarly, LCO has lower mean AUROC than LTO (p=0.015). In regression tasks the difference in PCC between LCO and LPO is significant (p=0.018). Due to the small sample sizes none of the other differences are statistically significant. Figure 1e shows the reported AUROC values per dataset. The distributions are relatively similar and no statistically significant differences can be shown between the datasets.